OpenStudy (anonymous):

Look at the coordinates to the at the bottom. What is the slope-intercept form of the equation of the line shown in the graph? ((3,0)this one is farther to the left),(this one is farther to the right(0-4) )

OpenStudy (anonymous):

slope is \[\frac{y_2-y_1}{x_2-x_1}=\frac{-4-0}{0-3}=\frac{4}{3}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what if it was the oppostie on the for the coordinates?

OpenStudy (anonymous):

not sure what you mean. the coordinates do not come "first" and "second" it is your choice

OpenStudy (anonymous):

i mean as in the one farther to the left is farther to right etc

OpenStudy (anonymous):

makes no difference which one you choose first as long as you are consistent

OpenStudy (anonymous):

example \[(2,-3), (4,7)\] you can do \[\frac{7-(-3)}{4-2}=\frac{10}{2}=5\] or you can do \[\frac{-3-7}{2-4}=\frac{-10}{-2}=5\] same thing either way

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

just don't mix them up and you will be fine

OpenStudy (anonymous):

thanks